# LaMET’s Asymptotic Extrapolation vs. Inverse Problem

### v1

@@ III.2 Inverse Problem
Inverse problems arise when there is a limited set of data points, yet there is an extraordinarily large number of parameters/degrees of freedom due to the absence of a coherent physical guidance. To proceed, one typically introduces specific parametrizations or models to reduce the degrees of freedom and make the fit tractable. However, if there are a large number of models possible, it is very difficult to assess the systematic errors, because it is hard to argue, based on physics, which model is better. Inverse problems are generally ill-posed, in contrast to well-posed problems. Mathematically, the ill-posedness of an inverse problem is characterized by three criteria proposed by Jacques Hadamard: existence, uniqueness, and stability of the solution with respect to the input data [215, 216]. A problem is considered ill-posed if it fails to meet any of these conditions. The ill-posedness typically manifests as an instability of the solution, which makes error estimation very difficult.
The global analysis of PDF from experimental data is a standard IP. As a continuous function of the momentum fraction $x$, PDFs are difficult to model. In fact, one might go so far as to say that a PDF is a collection of an infinite number of parameters. Of course, we believe that it is a continuous function and probably differentiable to a high degree except for $x=0$. A typical modelling of the PDF involves incorporating physics constraints like Regge behavior as $x\to 0$ and power counting in the $x\to 1$ limit. However, in the middle-$x$ region there is no physics-guided model, and one has to interpolate the two end-point behaviors. Thus, unless there is a large number of experimental data points with wide kinematic coverage (which, fortunately, we do have) and the continuity assumptions are imposed, IP algorithms tend to generate many solutions with hard-to-estimate systematic errors. Indeed, for small and large $x$, where data are sparse, the uncertainty of the PDF can become extremely large. It is challenging for IP methods to provide a reliable small-$x$ PDF in the absence of data constraints from very high-energy experiments.
The lattice QCD data analysis in the SDF framework is of a similar nature where one attempts to reconstruct a continuous PDF from limited segment of leading-twist correlation. A particular IP method within the SDF framework is introduced in Ref. [188], using the Gaussian Process Regression (GPR) with a logRBF kernel, defined as
$K\left(x,x^{\prime}\right)=\sigma^{2}\exp\left(-\frac{\left(\ln(x)-\ln\left(x^{\prime}\right)\right)^{2}}{2\ell^{2}}\right)\leavevmode\nobreak\ ,$ (6)
@@
Figure 2: An example of an inverse problem in the SDF framework: reconstruction of the PDF from short-distance correlation functions at $z=0.26$ fm with proton momenta $P^{z}=1.6,2.0,2.4,2.8,3.2$ GeV on the N203 ensemble, as reported in Ref. [89]. The reconstruction is performed using the GPR method with the logRBF kernel proposed in Ref. [188]. To illustrate the model uncertainty arising from different hyperparameter choices, two sets of priors are compared: the red band corresponds to $l=0.7\approx\ln(2)$, $\sigma^{2}=16$ and prior mean $f(x)=4$; the blue band corresponds to $l=0.2\approx\ln(1.25)$, $\sigma^{2}=2$ and prior mean $f(x)=1$.
With limited lattice data, IP methods that rely primarily on mathematical constraints and minimal physics input often lead to uncontrolled and unrealistic uncertainties. Unlike asymptotic fits, which use rigid, physically motivated functional forms with a small number of parameters, the IP method presented in Ref. [188] allows a broad space of solutions constrained only by abstract assumptions. While such flexibility is valuable for general problems, it limits reliability in extrapolating lattice data to the asymptotic region, where physical insight is essential. Until all the physical constraints in LaMET are fully integrated into IP frameworks, asymptotic analysis remains the most reliable approach for extrapolation.
@@ III.3 Asymptotic Extrapolation in LaMET
An example is the zero-momentum lattice correlation function for the nucleon transversity quasi-PDF shown in Fig. 3 from Ref. [89] [^1] Since the transversity case is a state-of-the-art LaMET analysis for nucleon PDFs, we will use it as our example throughout this comment paper., where the exponential decay is clearly visible in the data. We use the standard ground-state fitting to find the decay constant $m_{\rm eff}$ in Fig. 3(b). The plateau or asymptotic region is reached after $z\sim 0.75$ fm for all lattice ensembles, where the data still have relatively high precision. The mass obtained here corresponds to a ‘heavy-light” meson, which has a linear divergence from the “heavy quark” [65]. The linear divergence can be renormalized through a particular mass subtraction scheme (a constant shift in Fig. 3(b)), and the residual mass in the renormalized correlator can be chosen to be on the order of $\Lambda_{\rm QCD}$, say, $0.2-0.3$ GeV, which is correlated with the renormalon regulation scheme in the matching coefficient [207]. Note that this physical mechanism contradicts the interpretation in Ref. [1], which attributes the exponential decay to the pion mass.
@@
Figure 3: An example of high precision data in which we see clearly asymptotic exponential decay around $z=1$ fm.
@@
where $A(P^{z})$ denotes the product of pion-to–heavy-light meson transition form factors, which decrease with increasing $P^{z}$, reflecting a Lorentz contraction effect. Therefore, at nonzero $P^{z}$, the phase factor $e^{-iP^{z}z}$ predicts that the matrix element will change sign in the asymptotic region, a feature observed in many calculations, such as Refs. [1, 89], which can be incorporated into extrapolation models. A more accurate and general asymptotic form for extrapolation will be provided in [208].
While asymptotic extrapolation might be perceived as a model, it is in fact grounded in a rigorous derivation from the dispersion relation, effectively functioning as an EFT expansion. In the asymptotic limit, the allowed functional forms for lattice correlation functions are restricted to exponential decay, power-law behavior, and phase modulation for 4D Euclidean field theories with a mass gap. One typically begins with the leading behavior characterized by a finite number of parameters, and can systematically improve accuracy by including subleading terms.
If the lattice data within the sub-asymptotic region $[z_{SA},z_{A}]$ (for the nucleon, $z_{SA}\sim 0.5$ fm and $z_{A}\sim 1.2$ fm) have sufficiently small errors, then the extrapolation can be tightly constrained by the imposed physical conditions. This level of precision has been demonstrated for the renormalized transversity quasi-PDF matrix elements in Ref. [89], as shown in Fig. 4(a). A reliable extrapolation can be obtained by fitting the asymptotic model to data within a chosen window inside $[z_{SA},z_{A}]$. The truncation point $z_{L}$—beyond which data are replaced by the extrapolated form with its associated error band—may lie either below or above $z_{A}$, depending on the data precision.
@@
    The lattice data should be truncated at a point within the sub-asymptotic region where they have decreased smoothly to near zero.
@@
    The extrapolated results should continue to decrease in amplitude even with the presence of oscillations, consistent with asymptotic decay.
The above strategy has been used in most LaMET analyses so far. When all conditions are strictly met, the FT errors are well under control at finite $x$ and typically smaller than the other systematic uncertainties.
@@
The claim about the inverse problem in LaMET in Ref. [1] focuses on the practical concern that noisy lattice data in the asymptotic region may prevent a reliable asymptotic extrapolation and, therefore, should be analyzed by IP methods with minimal physics constraints. The authors also provided numerical evidence through tests on a set of nucleon matrix elements, demonstrating that variations among different IP methods with exponential decay still result in considerable differences from moderate to large $x$. This is consistent with our expectation that IP methods with little physics input will generate large errors.
However, we still expect that the asymptotic extrapolation in LaMET is a better option for providing a reliable error estimate even when the data quality is not ideal. Besides, the arguments and data analysis presented in Ref. [1] are problematic. First, data precision in the asymptotic region can be improved with realistic resources and new lattice techniques [217, 218, 219, 70, 142], which the authors described as exponentially hard. In addition, the numerical tests in Ref. [1] suffer from several problems, including large errors in the asymptotic region, an incorrect renormalization scheme, and analysis methods that do not satisfy physical constraints. Ultimately, the uncertainty of the FT within the models remains bounded by our estimate from asymptotic extrapolation, which can serve as a realistic quantification of the error given the noisy data. Therefore, FT in fact remains under control in the standard LaMET analysis, contrary to the conclusion drawn in [1].
@@ IV.2 Data Quality and Renormalization Problems
Before commenting on the data quality, we would like to point out that using the ratio scheme [144] at large $z$ to obtain the quasi-PDFs is incorrect. As discussed in Sec. III, we expect the non-zero momentum matrix element to decay exponentially at the same rate as that in the rest frame. In the ratio scheme, the exponential decay cancels, which completely distorts the non-perturbative physics in the asymptotic region. However, the matrix element could still approach zero due to the remaining dependence on $z$ or $\lambda$, as suggested by Eq. (7). A serious LaMET analysis should have been aborted at this point. In order to proceed, we will base our comments on this “pseudo data.”
As shown in Figure 3 of Ref. [1], the ratio of matrix elements changes sign and is consistent with zero with increasingly large error bars in the range $118<\lambda<$ 11 (or $0.75\leavevmode\nobreak\ \text{fm}<z<1.03\leavevmode\nobreak\ \text{fm}$). This ratio after the first zero is small in amplitude due to the Lorentz contraction effect in Eq. (7). Therefore, most of the conclusions in Ref. [1] were drawn by analyzing this imprecise ratio, which should be insignificant to the FT when exponential decay is preserved with proper renormalization, such as in the hybrid scheme [84, 76, 207]. However, the sizable errors of the data can lead to significant variation in the fits and extrapolations, contributing to the final uncertainty of the FT.
This situation, if interpreted as pseudo data within the LaMET framework, can be significantly improved by reducing the errors in the region $0.75\leavevmode\nobreak\ \text{fm}<z<1.03\leavevmode\nobreak\ \text{fm}$ by a factor of two or more, allowing the truncation point to be set at $z_{L}\sim 1.03$ fm to better satisfy the physical conditions for asymptotic extrapolation. This precision is achievable with a four-fold increase in statistics or by using the kinematically enhanced interpolation operator [217], both of which are feasible nowadays.
@@ IV.3 Data-Driven IP Methods May Conflict with Physical Constraints
If one treats the ratio-scheme matrix elements as pseudo data for LaMET, then the first issue is exemplified by the Gaussian-decay model used in Ref. [1]. Although it is known to violate the laws of physics, as the nucleon correlation functions decay exponentially at long range  [^2] In fact, the exponential decay cancels out in the ratio scheme, but it is even more difficult to justify why the Gaussian-decay would survive., the punishment here is not significant.
@@
Unlike the “Exact exp” approach, which follows the LaMET procedure of truncation and extrapolation, the GPR methods infer posterior distributions based on all data from $z=0$ to $1.13$ fm, despite the fact that exponential decay does not apply at small or moderate $z$. Moreover, despite being motivated by somewhat vague arguments listed in Ref. [188], the choice of RBF kernels and their associated hyperparameters in the GPR framework remains ad hoc and subject to model dependence [1, 173, 188], and there is no clear mechanism to ensure that they do not distort the physical conditions in the process. As we show below, this can result in unphysical extrapolations—even with high-precision data. Finally, the indiscriminate use of all noisy data also unnecessarily amplifies the fitting errors.
@@
After all, to further analyze the issue with the GPR fits using RBF-based kernels, we will temporarily treat them as comparable to the asymptotic extrapolation. As it turns out, none of them satisfies all the physical conditions described in Sec. III.3. As shown in Figures 3 and 4 in Ref. [1], the RBF-exp and logRBF-exp fits match the data up to $\lambda\sim 8$ (or $z\sim 0.75$ fm), beyond which they begin to deviate, giving the appearance of extrapolations starting at $\lambda_{L}\sim 8$. However, the matrix elements and their errors in the extrapolation region continue to grow in magnitude before slowly decaying as $\lambda\to\infty$, clearly violating the third physical condition. In particular, even the “Exact exp” extrapolation slightly violates the first physical condition, since it truncates at $z\sim 0.75$ fm, which is early for the amplitude of $h(z,P)$—including its slope—to decay sufficiently close to zero.
@@
For the purpose of comparison, we conducted analogous tests employing identical hyperparameters, as indicated in Reference 52 of Ref. [1], using the hybrid-scheme nucleon transversity quasi-PDF matrix elements from Ref. [89]. As shown in Fig. 4(a), these matrix elements are much more precise than those in Ref. [1] in the region $118<\lambda<$ 11 (or $0.85\leavevmode\nobreak\ \text{fm}<z<1.19\leavevmode\nobreak\ \text{fm}$), which allows for truncation at $z_{L}=1.19$ fm for asymptotic extrapolation, thus significantly reducing the uncertainty of the FT. However, despite using the same range of precise data in $z\leq z_{L}$, the GPR methods produce posteriors whose amplitudes and errors far exceed those of all data in the range $116<\lambda<$ 11, clearly violating the third physical condition. This suggests that the GPR methods used in Ref. [1] did not incorporate the exponential decay behavior in their prior settings, particularly when improper correlation lengths are chosen, causing the long-range behavior to be overly influenced by the small-$z$ region.
(a)  Nucleon transversity quasi-PDF matrix elements in the hybrid scheme at $P^{z}=1.82$ GeV from a lattice of spacing $a=0.085$ fm [89]. The Exact exp, RBF-exp, and logRBF-exp models with different correlation lengths $r$ [1] are used to extrapolate from $z_{L}=14a=1.19$ fm. The red data points are fitted exclusively by the Exact exp model..
@@
Therefore, the best approach to extrapolating the noisy pseudo data in Ref. [1] is through the asymptotic analysis. Its uncertainty can be better controlled by using the decay constant $m_{\rm eff}$—fitted from the zero-momentum matrix element—as input, truncating at a larger $z_{L}$, and adopting a more accurate asymptotic form for both extrapolation and error estimation.

### v2

@@ III.2 Inverse Problem
IPs arise when there is a limited set of data points, yet there is an extraordinarily large number of parameters or degrees of freedom due to the absence of a coherent physical guidance. To proceed, one typically introduces specific parametrizations or models to reduce the degrees of freedom and make the fit tractable. However, if there are a large number of models possible, it is very difficult to assess the systematic errors, because it is hard to argue, based on physics, which model is better. IPs are generally ill-posed, in contrast to well-posed problems. Mathematically, the ill-posedness of an IP is characterized by three criteria proposed by Jacques Hadamard: “existence, uniqueness, and stability of the solution with respect to the input data” [249, 250]. A problem is considered ill-posed if it fails to meet any of these conditions. The ill-posedness typically manifests as an instability of the solution, which makes error estimation very difficult.
The global analysis of PDF from experimental data is a standard IP. As a continuous function of the momentum fraction $x$, PDFs are difficult to model. In fact, one might go so far as to say that a PDF is a collection of an infinite number of parameters. Of course, we believe that it is a continuous function and probably differentiable to a high degree except for $x=0$. A typical modelling of the PDF involves incorporating physics constraints like Regge behavior as $x\to 0$ and power counting in the $x\to 1$ limit. However, in the middle-$x$ region there is no physics-guided model, one has to interpolate the two end-point behaviors. Thus, unless there is a large number of experimental data points with wide kinematic coverage—which we fortunately have—and the smoothness assumptions are imposed, IP algorithms tend to generate many solutions with hard-to-estimate systematic errors. Indeed, at small and large $x$ where the data are sparse, the uncertainty of the PDF can become extremely large. It is challenging for IP methods to provide a reliable small-$x$ PDF in the absence of data constraints from high-energy experiments.
The lattice QCD data analysis in the SDF framework is of a similar nature where one attempts to reconstruct a continuous PDF from a limited segment of leading-twist correlation. The model uncertainty in fitting the PDFs is difficult to quantify and has been shown to be significantly underestimated when using rigid parameterizations of their functional form—for example, see Refs. [251, 252]. A particular IP method within the SDF framework was introduced in Ref. [224], using the Gaussian Process Regression (GPR) with a logRBF kernel, defined as
$K\left(x,x^{\prime}\right)=\sigma^{2}\exp\left(-\frac{\left(\ln(x)-\ln\left(x^{\prime}\right)\right)^{2}}{2\ell^{2}}\right)\penalty 10000\ ,$ (9)
@@
Figure 2: An example of an IP in the SDF framework: reconstruction of the PDF from short-distance correlation functions at $z=0.26$ fm with proton momenta $P^{z}=1.6,2.0,2.4,2.8,3.2$ GeV on the N203 ensemble, as reported in Ref. [111]. The reconstruction is performed using the GPR method with the logRBF kernel proposed in Ref. [224]. To illustrate the model uncertainty arising from different hyperparameter choices, two sets of priors are compared: the red band corresponds to $l=0.7\approx\ln(2)$, $\sigma^{2}=16$ and prior mean $f(x)=4$; the blue band corresponds to $l=0.2\approx\ln(1.25)$, $\sigma^{2}=2$ and prior mean $f(x)=1$.
With limited lattice data, IP methods that rely primarily on mathematical constraints and minimal physics input often lead to uncontrolled and unrealistic uncertainties. Unlike asymptotic fits, which use rigorous, physically motivated functional forms with a small number of parameters, the IP method presented in Ref. [224] allows a broad space of solutions constrained only by abstract assumptions. While such flexibility is valuable for general problems, it limits reliability in extrapolating lattice data to the asymptotic region, where physical insight is essential. Until all the physical constraints in LaMET are fully integrated into IP frameworks, asymptotic analysis remains the most reliable approach for extrapolation.
@@ III.3 Asymptotic Extrapolation in LaMET
An example is the zero-momentum lattice correlation function for the nucleon transversity quasi-PDF shown in Fig. 3 from Ref. [111] [^1] Since Ref. [111] is a state-of-the-art LaMET analysis for nucleon PDFs, we will use it as our example throughout this paper., where the exponential decay is clearly visible in the data. We use the standard ground-state fitting to find the decay constant $m_{\rm eff}$ in Fig. 3(b). The plateau or asymptotic region is reached after $z\sim 0.75$ fm for all lattice ensembles, where the data still have relatively high precision. The mass obtained here corresponds to a “heavy-light” meson, which has a linear divergence from the “heavy quark” [97]. The linear divergence can be renormalized through a particular mass subtraction scheme (a constant shift in Fig. 3(b)), and the residual mass in the renormalized correlator can be chosen to be on the order of $\Lambda_{\rm QCD}$, say, 0.20.30.2-0.3 GeV, which depends on the renormalon regularization scheme in the matching coefficient [241]. Since the light-cone PDF is renormalon free, this scheme dependence is canceled by the corresponding matching coefficient. Therefore, one can choose a proper scheme to shift the Wilson-line mass subtraction within $O(\Lambda_{\rm QCD})$ while maintaining the perturbative convergence and numerical stability of matching, so that the renormalized correlation function decays faster. This does not change the final uncertainty in the PDF but makes the FT easier to control. Note that this physical mechanism contradicts the interpretation in Ref. [1], which attributes the exponential decay to the pion mass.
@@
Figure 3: An example of high-precision data in which we see clearly asymptotic exponential decay around $z=1$ fm.
@@
where $A(P^{z})$ denotes the product of pion-to-heavy-light meson transition form factors, which decrease with increasing $P^{z}$, reflecting a Lorentz contraction effect. Therefore, at nonzero $P^{z}$, the phase factor $e^{-iP^{z}z}$ predicts that the matrix element will change sign in the asymptotic region, a feature observed in many calculations, such as Refs. [1, 111], which can be incorporated into extrapolation models.
While asymptotic extrapolation might be perceived as a model, it is in fact grounded in a rigorous derivation from the dispersion relation, effectively functioning as an EFT expansion. In the asymptotic limit, one typically begins with the leading behavior characterized by a finite number of parameters, and can systematically improve accuracy by including subleading terms. A more accurate and general asymptotic form for extrapolation will be provided in [242].
Of course, understanding the long tail of correlation functions—such as excited-state contamination—remains a challenging problem in many scenarios. For example, in the hadronic vacuum polarization (HVP) contribution to the muon’s anomalous magnetic moment, control over the tail is achieved through spectral decomposition [253]; by exploiting the positivity of the correlator to impose rigorous upper and lower bounds [254, 255, 256, 257, 258]; or by using Bayesian model averaging to systematically combine multiple fit models and account for model-selection uncertainty [259, 260, 261]. The asymptotic extrapolation in LaMET faces similar issues and could benefit from techniques developed in such settings with EFT constraints. However, unlike the HVP case—which involves integrating the correlation function with a kernel that enhances the long-range contribution—the FT kernel $\exp(ixP^{z}z)$ in LaMET naturally suppresses this contribution at finite $x$. Even if uncertainty remains in reconstructing the long tail, the FT kernel still ensures an upper bound on it, provided that the exponential decay is encoded. We demonstrate this in the example in Sec. IV.4 below.
If the lattice data within the sub-asymptotic region $z_{SA},z_{A}]$ (for the nucleon, $z_{SA}\sim 0.5$ fm and $z_{A}\sim 1.2$ fm) have sufficiently small errors, then the asymptotic extrapolation can be tightly constrained. This level of precision has been demonstrated for the renormalized transversity quasi-PDF matrix elements in Ref. [111], as shown in Fig. 4(a). A reliable extrapolation can be obtained by fitting the asymptotic form to data within a chosen window inside $z_{SA},z_{A}]$. The truncation point $z_{L}$—beyond which data are replaced by the extrapolated form with its associated error band—may lie either below or above $z_{A}$, depending on the data precision.
@@
    The lattice data should be truncated at a point within the sub-asymptotic and asymptotic regions where they have decreased smoothly to near zero.
@@
    The extrapolated result—including its error—should continue to decrease with increasing $z$, consistent with the asymptotic decay. In cases where oscillations or sign changes are present, the amplitude of oscillation should decay exponentially.
Since the exact asymptotic behavior of a correlation function depends on the operator and external state, the parameters used in the extrapolation must be determined on a case-by-case basis. After performing the extrapolation, it remains necessary to estimate the model uncertainty in the FT. While most LaMET analyses have imposed the physical conditions for extrapolation, not all have provided an estimate of the associated uncertainty, as criticized by Ref. [1]. In Ref. [97], such an estimate was obtained by varying the truncation point  [^2] In Ref. [240], it was pointed out that the pion data points beyond $z_{L}\sim 0.9$ fm should not be discarded, as they are statistically nonzero. However, since their magnitude continues to grow to unphysically large values, they were ruled out in Ref. [97] based on the pion’s size, such as its charge radius of about $0.6$ fm [262]. After all, this is a special case of LaMET analyses. In the future, such a restriction will be relaxed with more dedicated lattice calculations of the long-range correlations., the prior for the decay constant, and the extrapolation model. Notably, unphysical power-law decay models were also included to emphasize the differences. It was found that the variations among extrapolations are negligible at moderate $x$, so they are not included in the final error budget.
Nevertheless, when the precision of data in the sub-asymptotic region is not sufficiently high, the FT uncertainty can become non-negligible. In such cases, the method in Ref. [97] may be less reliable, as it does not exhaust the model space. For a more accurate quantification of this error, one can first determine $m_{\rm eff}$, $z_{SA}$ and $z_{A}$ from high-precision zero-momentum matrix elements and then use them as priors at nonzero momenta. Subsequently, techniques analogous to those used in HVP calculations can be applied to constrain the long-distance tail and yield a more reliable uncertainty estimate [242]. Since the asymptotic decay arises from an EFT expansion in which subleading terms fall off more rapidly than the leading contribution, this uncertainty is controllable. A more accurate derivation of the asymptotic form and its corrections [242] will refine this estimate.
@@
The claim about the IP in LaMET in Ref. [1] focuses on the practical concern that noisy lattice data in the sub-asymptotic region may prevent a reliable asymptotic extrapolation and, therefore, should be analyzed by IP methods with minimal physics constraints. The authors also provided numerical evidence through tests on a set of nucleon matrix elements, demonstrating that variations among different IP methods with exponential decay still result in considerable differences from moderate to large $x$. This is consistent with our expectation that IP methods with minimal physics input will lead to unnecessarily conservative errors.
However, we still expect that the asymptotic extrapolation in LaMET is a better option for providing a reliable error estimate even when the data quality is not ideal. Besides, the arguments and data analysis presented in Ref. [1] are problematic. First, data precision in the asymptotic region can be improved with realistic resources and new lattice techniques [263, 264, 265, 126, 133], which the authors described as exponentially hard. In addition, the numerical tests in Ref. [1] suffer from several problems, including large errors in the sub-asymptotic region, an incorrect renormalization scheme, and analysis methods that do not satisfy physical constraints. Ultimately, the uncertainty of the FT within the models remains bounded by our estimate from asymptotic extrapolation, which can serve as a realistic quantification of the error given the data quality. Therefore, FT in fact remains under control in the standard LaMET analysis, contrary to the conclusion drawn in [1].
@@ IV.2 Data Quality and Renormalization Problems
Before commenting on the data quality, we would like to point out that using the ratio scheme [173] at large $z$ to obtain the quasi-PDFs is incorrect. First, it introduces uncontrolled non-perturbative effects at intermediate and large $z$, which subsequently contaminate all $x$ after the FT. Second, there is no QCD factorization for a ratio-scheme quasi-PDF in $x$-space despite the studies in early literature [175, 163, 45]. In fact, $x$-space factorization has been rigorously established for the $\overline{\mathrm{MS}}$ scheme only, implying that any scheme not perturbatively convertible to it at all zz—such as the ratio scheme—is not factorizable.
Besides, since Ref. [1] aims to demonstrate a problem in LaMET analysis, it should at least adhere to the same procedure when making this claim. As discussed in Sec. III, we expect the non-zero momentum matrix element to decay exponentially at the same rate as that in the rest frame. Taking a ratio of matrix elements at different momenta removes this key feature entirely, so no definitive conclusion can be drawn from such an example. Hence, the ratio scheme is logically inconsistent for the purpose of Ref. [1]. In order to proceed, we will treat the ratio-scheme matrix elements as “pseudo data” for LaMET analysis.
As shown in Figure 3 of Ref. [1], the ratio of matrix elements changes sign and is consistent with zero with increasingly large error bars in the range $8<\lambda<11$ (or $0.75\penalty 10000\ \text{fm}<z<1.03\penalty 10000\ \text{fm}$). This ratio after the first zero is small in amplitude due to the Lorentz contraction effect in Eq. (10). Most of the conclusions in Ref. [1] were drawn by analyzing this imprecise ratio, which should be insignificant to the FT when exponential decay is preserved with proper renormalization such as the hybrid scheme [78, 128, 241]. However, the sizable errors of data can lead to significant variation in the fits and extrapolations, contributing to the uncertainty of the FT.
This situation, if interpreted as pseudo data within the LaMET framework, can be significantly improved by reducing the errors in the region $0.75\penalty 10000\ \text{fm}<z<1.03\penalty 10000\ \text{fm}$ by a factor of two or more, allowing the truncation point to be set at $z_{L}\sim 1.03$ fm to better satisfy the physical conditions for asymptotic extrapolation. This precision is achievable with a four-fold increase in statistics or by using the kinematically enhanced interpolation operator [263], both of which are feasible nowadays.
@@ IV.3 Data-Driven IP Methods May Conflict with Physical Constraints
If one treats the ratio-scheme matrix elements as pseudo data for LaMET, then the first issue is exemplified by the Gaussian-decay model used in Ref. [1]. Although it is known to contradict the exponential decay of nucleon correlation functions at long range  [^3] In fact, the exponential decay cancels out in the ratio scheme, but it is even more difficult to justify why the Gaussian-decay would survive., the punishment here is not significant.
@@
Unlike the Exact exp approach, which follows the LaMET procedure of truncation and extrapolation, the GPR methods infer posterior distributions based on all data from $z=0$ to $1.13$ fm, despite the fact that exponential decay does not apply at smaller $z$. Moreover, despite being motivated by somewhat vague arguments listed in Ref. [224], the choice of RBF kernels and their associated hyperparameters in the GPR framework remains ad hoc and subject to model dependence [1, 183, 224], and there is no clear mechanism to ensure that they do not distort the physical conditions in the process. As we show below, this can result in unphysical extrapolations—even with high-precision data. Finally, the indiscriminate use of all noisy data also unnecessarily amplifies the fitting errors.
@@
After all, to further analyze the issue with the GPR fits using RBF-based kernels, we will temporarily treat them as comparable to the asymptotic extrapolation. As it turns out, none of them satisfies all the physical conditions described in Sec. III.3. As shown in Figures 3 and 4 in Ref. [1], the RBF-exp and logRBF-exp fits match the data up to $\lambda\sim 8$ (or $z\sim 0.75$ fm), beyond which they begin to deviate, giving the appearance of extrapolations starting at $\lambda_{L}\sim 8$. However, the matrix elements and their errors in the extrapolation region continue to grow in magnitude before slowly decaying as $\lambda\to\infty$, clearly violating the third physical condition. In particular, even the Exact exp extrapolation slightly violates the first physical condition, since it truncates at $z\sim 0.75$ fm, which is early for the amplitude of h(z,P)h(z,P)—including its slope—to decay sufficiently close to zero.
@@
For the purpose of comparison, we conducted analogous tests employing identical hyperparameters  [^4] Unlike Ref. [1], we do not impose a threshold on the uncertainty of the reconstructed tail when determining the hyperparameters, because such a choice is arbitrary. Our purpose is to show that GPR without physical constraints could lead to unphysical extrapolations. While adopting a threshold might help the GPR reconstruction satisfy the conditions in Sec. III.3, its physical basis is unclear, so the associated model uncertainty cannot be reliably quantified., as indicated in Reference 52 of Ref. [1], using the hybrid-scheme nucleon transversity quasi-PDF matrix elements from Ref. [111]. As shown in Fig. 4(a), these matrix elements are much more precise than those in Ref. [1] in the region $8<\lambda<11$ (or $0.85\penalty 10000\ \text{fm}<z<1.19\penalty 10000\ \text{fm}$), which allows for truncation at $z_{L}=1.19$ fm for asymptotic extrapolation, thus significantly reducing the uncertainty of the FT. However, despite using the same range of precise data in $z\leq z_{L}$, the GPR methods produce posteriors whose amplitudes and errors far exceed those of all data in the range $6<\lambda<11$, clearly violating the third physical condition. This suggests that the GPR methods used in Ref. [1] did not incorporate the exponential decay behavior in their prior settings, particularly when improper correlation lengths are chosen, causing the long-range behavior to be overly influenced by the smaller-$z$ region.
(a) Nucleon transversity quasi-PDF matrix elements in the hybrid scheme at $P^{z}=1.82$ GeV from a lattice of spacing $a=0.085$ fm [111]. The Exact exp, RBF-exp, and logRBF-exp models with different correlation lengths $r$ [1] are used to extrapolate from $z_{L}=14a=1.19$ fm. The red data points are fitted exclusively by the Exact exp model.
@@
Therefore, the best approach to extrapolating the pseudo data in Ref. [1] is through the asymptotic analysis. Its uncertainty can be better controlled by using the decay constant $m_{\rm eff}$—fitted from the zero-momentum matrix element—as input, truncating at a larger $z_{L}$, and adopting a more accurate asymptotic form for both extrapolation and error estimation.